OpenStudy (anonymous):

Why is y(t) = x^2(t) not linear? rule: ax1(t) + bx2(t) -> ay1(t) + by2(t)

OpenStudy (anonymous):

Just to clarify, is that (x^2)*(t) or x^(2t)?

OpenStudy (anonymous):

its (x^2)(t)

OpenStudy (anonymous):

Alright let me type this out, one sec...

OpenStudy (anonymous):

oke thanks :)

OpenStudy (anonymous):

if i use the rule: a(x1^2)(t) + b(x2^2)(t) -> ay1(t) + by2(t) BUT! ay1(t) = a(x1^2)(t) ay2(t) = a(x2^2)(t) so its linear?! but ISNT.... am i doing something wrong?

OpenStudy (anonymous):

Yeah i'll explain it. its just taking a bit to type.

OpenStudy (anonymous):

okeee... thats why i am asking this question :) i will wait :)

myininaya (myininaya):

\[y(t)=x^2(t)\] \[ax_1(t) + bx_2(t) -> ay_1(t) + by_2(t)\] \[ax_1(t)+bx_2(t)->ax_1^2(t)+bx_2^2(t)\] its not linear because this is not true \[x_1^2 \neq x_1 \text{ and } x_2^2 \neq x_2\]

OpenStudy (anonymous):

So what I would do on this is take the derivative with respect to t. The derivative gives you the equation for the slope of the original equation. In order for an equation to be considered linear, its slope must be linear. So for this equation, when you take the derivative with respect to t, you get: \[\left(\begin{matrix}d \\ dt\end{matrix}\right)[y(t)=(x^2)(t)]\rightarrow y'(t)=x^2\] A linear slope would be something like y'(t)=x or y'(t)=4x. However the slope of this line is y'(t)=x^2 which is quadratic, so the initial equation as a whole would be considered quadratic. Make sense??

OpenStudy (anonymous):

The answer before mine also works but I just like this way better

OpenStudy (anonymous):

wait guys... i have to use the rule ive mentioned

OpenStudy (anonymous):

Alright well in that case, use the other person's answer. I could explain the other answer too if you don't get it

OpenStudy (anonymous):

@myininaya @ jpomer325 you are saying that \[x1 ^{2}\neq x1\] but i have to multiply my x1^2 with a but why is it still x1?

myininaya (myininaya):

jpomer i'm really confused about your answer why didn't you differentiate the other side and y'=constant implies y is linear

OpenStudy (anonymous):

i dont get it.. i already wrote what i did.. but whats wrong with my solution?

OpenStudy (anonymous):

i dont have to differentiate anything... i only have to use the rule from above.. but i dont get the solution of jpomer325 :s

OpenStudy (anonymous):

haha its alright i didnt use those rules anyway. a and b are just constants so you can ingnore them for all practical purposes in this problem

OpenStudy (anonymous):

myininaya i really hope you can help me with this question :)

OpenStudy (anonymous):

\[y(t) = x ^{2}(t)\] \[\alpha(x1^{2})(t) + \beta(x2^{2})(t) = \alpha(y1^{2})(t) +\beta(y2^{2})(t) \]

OpenStudy (anonymous):

\[y(t)=(x^2)(t)\] Using the above equation, she plugged in (x^2)(t) for all y(t)'s in this equation: \[ax_1(t)+bx_2(t)−>ay_1(t)+by_2(t)\] Resulting in the following: \[ax_1(t)+bx_2(t)−>a(x_1)^{2}(t)+b(x_2)^{2}(t)\] However, this statement is obviously not true, because by definition: \[x _{1}\neq(x _{1})^{2}\] and \[x _{2}\neq(x _{2})^{2}\] So the rule for linearity is broken by the given equation y(t)=(x^2)(t).

OpenStudy (anonymous):

but why dont i have to apply ax1(t) as a(x1^2)(t)?

OpenStudy (anonymous):

Well that is what you are doing, comparing them to see if the statement is true. But for simplicity's sake, you can leave out the coefficients "a" and "b" because they just cancel out. \[[a(x)=a(x^2)]/a \rightarrow x=x^2\]

OpenStudy (anonymous):

Thats just an example of how it cancels

OpenStudy (anonymous):

okeee... but could you explain why y(t) = 2x(t) + 3 is also not linear?

OpenStudy (anonymous):

is there a rule with that one too or not

OpenStudy (anonymous):

the same rule has to be applied

OpenStudy (anonymous):

Alright well for this one, I don't see how its not linear either, because the quick check I was always taught was that as long as the y(t) and x(t) terms were linear (and not (x^3)(t) or (y^2)(t) for example), then the equation is linear.

OpenStudy (anonymous):

it is not linear :S i have to apply the same rule. but i dont get it at all -.-

OpenStudy (anonymous):

Well let me see what happens if we plug in the same way we did on the first problem...

OpenStudy (anonymous):

\[y(t)=2x(t)+3\] Using the above equation, we plug in 2x(t)+3 for all y(t)'s in this equation: \[ax_1(t)+bx_2(t)−>ay_1(t)+by_2(t)\] Resulting in the following: \[a(x_1)(t)+b(x_2)(t)−>a*(2x_1(t)+3)+b*(2x_2(t)+3)\] However, this statement is obviously not true, because by definition: \[x_1(t)≠2x_1(t)+3\] and \[x_2(t)≠2x_2(t)+3\] So the rule for linearity is broken by the given equation y(t)=2x(t)+3.

OpenStudy (anonymous):

Pretty much the same as the first one. But I'm also still not entirely sure thats all there is to it. Might be about that time to say f*ck it. You'll go crazy otherwise haha

OpenStudy (anonymous):

haha thats funny it changes f word to flutter

OpenStudy (anonymous):

hmm.. that makes sence.. but this one? y(t) = Ax(t) this one is linear

OpenStudy (anonymous):

my teachers said that i had to fill the formula also in for ax1(t) + bx2(t)

OpenStudy (anonymous):

Well since I would assume A is a coefficient, y(t) would vary linearly with x(t), so its linear

OpenStudy (anonymous):

Where did they say you plug that in?

OpenStudy (anonymous):

i dont know.. but i had to use this rule... and use it for these examples..

OpenStudy (anonymous):

Yeah I really dont know. I'd have to see the whole problem or the textbook its based out of

myininaya (myininaya):

whats the question?

OpenStudy (anonymous):

why y(t) = Ax(t) is linear

myininaya (myininaya):

use your definition for linearity

OpenStudy (anonymous):

and why are you only filling in for ay1(t) + by2(t) and not the ax1(t) + bx2(t)

OpenStudy (anonymous):

my teacher said that i also had to fill in the ax1(t) + bx2(t)

myininaya (myininaya):

\[y(t)=x^2(t)\] ok maybe you like this more: we have \[y_1(t)=x_1^2(t)\] \[y_2(t)=x_2^2(t)\] \[\text{ Let } x_3(t)=ax_1(t)+bx_2(t) \text{ then } y_3(t)=x_3^2(t)=(ax_1(t)+bx_2(t))^2\] \[=a^2x_1^2(t)+2abx_1(t)x_2(t)+b^2x^2_2(t)\] \[\text{ but } ay_1(t)+by_2(t) = ax_1^2(t)+bx_2^2(t)\]

jhonyy9 (jhonyy9):

so the anwser on your question is : why ... ? so because we know that every quadratic equation has graph parabola what not equal linear ...